# Efficient Adaptable High-throughput Streaming Aggregation Engines

### v3

@@ Abstract
Aggregation queries are a series of computationally-demanding analytics operations on counted, grouped or time series data. They include tasks such as summation or finding the median among the items of the same group, and within a specified number of the last observed tuples for sliding window aggregation (SWAG). They have a wide range of applications including database analytics, operating systems, bank security and medical sensors. Existing challenges include the hardware complexity that comes with efficiently handling per-group states using hash-based approaches. This paper presents an adaptable pipeline for calculating a wide range of aggregation queries with high throughput. It is then adapted for SWAG and achieves up to 476x speedup over the CPU core of the same platform. It achieves unparalleled levels of performance and functionality such as a throughput of 1 GT/s on our setup for SWAG without groups, and more advanced operators with up to 4x the window sizes than the state-of-the-art with groups as an approximation for SWAG featuring per-group windows using a fraction of the resources and no DRAM.
Index Terms:
FPGA, reconfigurable, group-by-aggregate, engine, prefix scan, sliding window, SWAG, key-value pairs, median
@@ I Introduction
Aggregation is the task of combining raw data to produce more useful and summarised knowledge. It appears in many disciplines, especially relating to distributed and parallel programming models, such as MapReduce [1], where multiple workers/nodes produce a partial result that needs aggregating. The worker notion is also popular in FPGA designs [2]. Equivalently, an established software design pattern is the fork-join model, which is a prominent concurrency pattern with processes or threads calculating partial results [3].
An emerging need in data analytics is the use of programming models and related infrastructure to be able to benefit from the increasingly-wide memories and datapaths [4]. A more traditional alternative is to have multiple workers and ports, as with FPGA-based database designs with multi-threading-equivalent arrangements. However, since the memory access pattern is highly-affected by this design choice, the hardware complexity is also affected, such as by co-designing cache or hash structures on the programmable logic. In the software world, this has already been reflected in the prominence of the key-value store paradigm, where the simplified storage model can enable higher throughput [5].
@@
A similar category of queries is sliding window aggregation (SWAG) [1]. A time domain is associated with the data, as with queries specifying a time range to form windows for aggregation. Due to the fact that the window formation is applied on-the-fly, near where the data is gathered, the stream is considered partly sorted with regard to accompanying timestamps or the implied order of the data. SWAG usually yields a random memory access pattern to calculate, including for fine-tuned CPU algorithms [6]. An example is when finding the median, as even with FPGAs, current designs require caching of larger states [7]. This complicates the design by increasing the on-chip and off-chip memory requirements for supporting/approaching the worst case (as with multiple inputs hashing to the same slot for hashsets).
Figure 1: Five-step pipeline of the proposed aggregation pipeline.
This paper presents a novel design for aggregation that integrates computation inside scalable structures such as the prefix scan and the butterfly network that process multiple elements per-cycle (high-throughput). The idea behind this design is the exploitation of those structures originally used for highly-optimised compaction [4] to simultaneously append aggregation information, leading to minimal resources (efficiency). The resulting base pipeline supports a plethora of operators and use cases (adaptable), as with the proposed adaptations for performing SWAG queries, by accompanying an on-chip sorter-cardinality counter. The resulting SWAG engine is highly memory-efficient to the point that it achieves similar functionality to highly-complex competing designs without their limiting requirement for an off-chip DRAM or compression. Conveniently, all logical states fit entirely inside on-chip memories such as block RAM (BRAM). For applications such as SWAG, the concept behind its advantageous system behaviour is the replacement of the need for large indexed or hashed structures by offloading the design complexity to sorting, as it eliminates fragmentation originally caused by maintaining per-group information. FPGA-based high-throughput sorters are thoroughly studied in literature [8, 9, 10, 11], but in order to seamlessly support the median operator in our SWAG engine, innovative lightweight algorithmic modifications are introduced.
The contributions of this paper are as follows:
@@ II Aggregation engine
The main idea behind the design is that the pre-existing prefix-scan topology of a compaction circuit can be exploited and reused to also calculate and append aggregation information to the data, before filtering redundant tuples. The focus of this pipeline is high-throughput, as it is always able to process up to $P$ elements per cycle. It is inspired by the PRRA architecture [4], where a rolling prefix scan is used to calculate the permutation indices for the permutation network (reverse butterfly) to achieve stream compaction (see figure 16, left).
@@ II-A Supported functionality
This design is a non-blocking building block working as a pipeline, i.e. it never asserts backpressure on the stream, as with the PRRA architecture. On the contrary, aggregation implies summarisation, hence the need to remove a newly generated intermediate sparsity with the compaction functionality of the permutation network [4]. For most aggregation tasks, this will result in only the last tuple per group being considered useful, since it is repurposed to carry all the relevant accumulated information, while the remaining aforementioned “redundant” tuples may carry partial results.
@@ II-B Pipeline steps
Figure 1 introduces the architecture of the proposed aggregation design, a high-throughput pipeline for calculating a wide range of aggregation tasks including group-by.
@@
the entities named $n$ are distributed inside the adapted prefix scan to calculate the permutation index of the tuples to achieve the round-robin effect by the reverse butterfly network, as with the PRRA [4]. At the same time they also operate on the keykey fields, which later becomes the resultresult field containing the aggregation summary per group. For example, if the query function is to find the sum of all the keys for each of the group IDs, the entities $n$ not only perform a cumulative sum for the reverse butterfly indexindex, but also for the key sum per group, until the last entry of the same group is processed to carry the full sum of the group. An equality check on the group field with the previous entry ($group_{i}\stackrel{{\scriptstyle?}}{{=}}group_{i-1}$) is enough for knowing where not to propagate the previous result. Using sum as an example, the start of a new groupgroup would leave the current keykey intact instead of adding the result of the last tuple.
@@
finally, the reverse butterfly network permutes the batch of results to achieve the round-robin effect in parallel (see figure 16, left), and is inline with [4]. The reverse butterfly network is used as an optimal-complexity switch, though other permutation networks can also be used. The lastlast flag is now renamed to a validvalid bit, as it now indicates that a port has a valid result.
@@ II-C Numerical example
Figure 3: High-throughput adaptable sliding window aggregation using a six-step pipeline.
@@
III Sliding window aggregation
The design from section II-B is adapted to perform sliding-window aggregation (SWAG) queries. In terms of functionality, the main difference from the base pipeline is that the input is processed in a windowed context.
@@ III Sliding window aggregation
SWAG queries are usually of the form “calculate the median of the last WS elements every WA tuples” [7], where median can be replaced by any supported aggregate operator, WS is the window size, and WA is the distance between the sampled windows. Figure 4 shows an instance of the average operator, where on every 2 tuples (WA) we find the average key within the last 4 tuples (WS). In the literature, WS is also referred to as the window range, and WA is also referred to as the bulk size, slide or stride. In some contexts, there is a group ID associated with each key, thus, the engine can also provide per group information, as with the base pipeline used for group-by aggregation.
@@ III-A Supported functionality
The proposed SWAG engine supports the operators of the base pipeline (see section II-A), plus selection-based operators including the median. Based on the query classification from [1], the supported types are Sum-like, Max-like and Sketch-like.
In contrast to software algorithms that calculate SWAG incrementally [1], which implement the window eviction post calculation [13], here, there is no requirement for an inverse function. This is because each window is processed in its entirety, which contributes to the adaptability of the SWAG engine for operators that cannot be calculated incrementally. Section VI elaborates on the applicability of the SWAG engine to other prominent aggregation workloads.
@@ III-B Design
The input data are seamlessly sorted on-the-fly using a small-scale sorter (see section III-D). As opposed to the base pipeline, the input is not expected to be sorted, since the latency of doing so is shown to be negligible for the window sizes of interest. Figure 3 introduces the pipeline of the SWAG engine in steps, for data arriving in batches of $P$ tuples.
@@
a sorter is needed right before the aggregation engine to sort the window based on the groupgroup IDs, as well as the keys. Note that sorting by the keys as well is not always required, such as when the operator is commutative as with summation. Since 4K elements, for example, are considered moderately large as SWAG window sizes [7], the sorting can entirely be done on-the-fly instead of indexing larger memories. FPGA-based high-throughput sorters are thoroughly studied in literature [8, 9, 10, 11].
@@
the last two steps reduce back to a PRRA architecture [4], as the remaining functionality only relates to the round-robin compaction. The final index calculation is done by the last prefix scan in the pipeline, while the first one is now responsible for all computation-related functionality including the other two cumulative sums.
@@ III-D Sorter with group-cardinality calculation
The ability to involve median calculation within a sorter depends on the sorter architecture [14]. Our proposed solution is based on a sorter that uses the linear sorter as a building block [15]. The linear sorter is the hardware equivalent of insertion sort, and is able to insert a new value into an already sorted sublist [16]. Each sorting cell, as depicted in figure 6 (center), has a comparator that compares its tuple with the incoming one. According to the result of the comparison and the one of its neighbour (“left smaller”), it either adopts the new value, propagates the left value, or remains as is.
@@
The baseline sorter of our design is able to provide high-throughput (multiple tuples per cycle) while supporting moderately large window sizes. The open-source sorter of [15] is specifically selected because it combines linear sorters with a merge tree (PMT [11]) for high-throughput. A special feature of this combination is that it is able to work efficiently in many-leaf (input list) merge mode using the same hardware, and this also applies to our new logic for resource efficiency.
@@
An additional modification is needed to propagate the correct group cardinalities, as summarised in figure 8. The parallel merge tree (PMT) needs to sum the cardinalities calculated by the linear sorters. Each FLiMS merger [17] of the PMT uses a network of compare-and-swap (CAS) units, which are sorters of 2-elements. The modification here is that each CAS unit uses the cond1 and cond2 conditions to determine how to update two incoming cardinalities every time they share the same group ID. The tuples carry two extra bits inside the mergers, one for denoting if the updating of their cardinalities has already happened, and one for their original source (lists A or B according to figure 8). Cond1 (to sum both cardinalities) is true whenever the cardinality update has not already happened to any of the 2 inputs, and the source is different. Cond2 (to adopt the other cardinality) is true whenever the cardinality update has happened to the other tuple, and the source is either different or the same. A similar arrangement is done for the selector stage units, as they are based on the same comparison, but they also remember any merged cardinalities from the elements of the last comparison. The latter is needed to support the case when a set of same group IDs is only coming from one of the lists A or B.
@@ IV Evaluation
This section evaluates the implementations of the group-by aggregate and SWAG engines as accelerators. All implementations are accompanied by a sorter/ cardinality counter of capacity $k$ of 128 elements, supporting arbitrary input sizes for the group-by engine by being applied recursively, and window sizes of up to $\textit{WS}=16K$ for the SWAG engine. The operator type is configured at run-time. The implemented engines support multiple operators also configurable at run-time, as summarised in table I.
@@ IV-A Experimental methodology
The engines are evaluated as co-processors in a Linux-running environment. They are implemented as AXI peripherals, and the codebase is mainly written in Verilog [^1] Source code (base): https://philippos.info/groupby/
Source code (SWAG): available soon
. The entities such as the compare-and-swap units for the sorter’s merger cells or the entities of the prefix scans are parameterisable, such as with custom data widths and functionality, and are written by hand. The structures that use those entities including the reverse butterfly network and the parallel merge tree in the sorter are generated by scripts. These are written in Python to easily infer such complex or repetitive entity topologies for arbitrary $P$ and $k$ (i.e. sorter size) values.
@@
The 2GB memory of the board is divided in two halves by the kernel; the kernel-mapped part, which is optimised for cached access to the CPU, and the non-kernel part, which is left for the FPGA. The target operating frequency of the AXI-related IPs is 250 MHz, with an internal frequency of 125 MHz for the adapted sorter and engines, complying with the sorter’s original design. This results in a 256-bit/cycle datapath for the engine from an 128-bit-wide HP port. This is relatively close to saturating the memory bandwidth of this platform [18]. Figure 9 visualises this experimental setup.
@@
The data first pass through the high-throughput sorter, and its output is directed towards the rest of the aggregation engine. The engine is either the group-by-aggregate version or the corresponding part of the SWAG engine. For the first, the sorter is the unmodified open-source IP from [15], whereas the latter combines the group cardinality functionality from section III-D.
@@
Their performance is compared against highly-optimised software, including reference implementations from IBM for SWAG [19]. The custom operators remain relatively simple, as they rely heavily on well-regarded functions and structures from C++’s standard template library (STL). The baseline software uses std::set and std::map to hold and update per-group information (notably similar performance to their hash-based unordered counterparts, not shown), and std::nth_element() to calculate the medians. The GCC compilation flags are -O3 -march=native.
IV-B Experimental results – Group-by aggregation

### v4

# Enthuse: Efficient Adaptable High-throughput Streaming Aggregation Engines

@@ Abstract
Aggregation queries are a series of computationally-demanding analytics operations on counted, grouped or time series data. They include tasks such as summation or finding the median among the items of the same group, and within a specified number of the last observed tuples for sliding window aggregation (SWAG). They have a wide range of applications including database analytics, operating systems, bank security and medical sensors. Existing challenges include the hardware complexity that comes with efficiently handling per-group states using hash-based approaches. This paper presents Enthuse, an adaptable pipeline for calculating a wide range of aggregation queries with high throughput. It is then adapted for SWAG and achieves up to 476x speedup over the CPU core of the same platform. It achieves unparalleled levels of performance and functionality such as a throughput of 1 GT/s on our setup for SWAG without groups, and more advanced operators with up to 4x the window sizes than the state-of-the-art with groups as an approximation for SWAG featuring per-group windows using a fraction of the resources and no DRAM.
@@ I Introduction
Aggregation is the task of combining raw data to produce more useful and summarised knowledge. It appears in many disciplines, especially relating to distributed and parallel programming models, such as MapReduce [Tangwongsan2018], where multiple workers/nodes produce a partial result that needs aggregating. The worker notion is also popular in FPGA designs [owaida2017centaur]. Equivalently, an established software design pattern is the fork-join model, which is a prominent concurrency pattern with processes or threads calculating partial results [de2014fork].
An emerging need in data analytics is the use of programming models and related infrastructure to be able to benefit from the increasingly-wide memories and datapaths [prra]. A more traditional alternative is to have multiple workers and ports, as with FPGA-based database designs with multi-threading-equivalent arrangements. However, since the memory access pattern is highly-affected by this design choice, the hardware complexity is also affected, such as by co-designing cache or hash structures on the programmable logic. In the software world, this has already been reflected in the prominence of the key-value store paradigm, where the simplified storage model can enable higher throughput [zhang2015mega].
@@
A similar category of queries is sliding window aggregation (SWAG) [Tangwongsan2018]. A time domain is associated with the data, as with queries specifying a time range to form windows for aggregation. Due to the fact that the window formation is applied on-the-fly, near where the data is gathered, the stream is considered partly sorted with regard to accompanying timestamps or the implied order of the data. SWAG usually yields a random memory access pattern to calculate, including for fine-tuned CPU algorithms [theodorakis2018hammer]. An example is when finding the median, as even with FPGAs, current designs require caching of larger states [geethakumari2023stream]. This complicates the design by increasing the on-chip and off-chip memory requirements for supporting/approaching the worst case (as with multiple inputs hashing to the same slot for hashsets).
Figure 1: Five-step pipeline of the proposed aggregation pipeline of Enthuse.
This paper presents Enthuse, a novel design for aggregation that integrates computation inside scalable structures such as the prefix scan and the butterfly network that process multiple elements per-cycle (high-throughput). The idea behind this design is the exploitation of those structures originally used for highly-optimised compaction [prra] to simultaneously append aggregation information, leading to minimal resources (efficiency). The resulting base pipeline supports a plethora of operators and use cases (adaptable), as with the proposed adaptations for performing SWAG queries, by accompanying an on-chip sorter-cardinality counter. The resulting SWAG engines are highly memory-efficient to the point that they achieves similar functionality to highly-complex competing designs without their limiting requirement for an off-chip DRAM or compression. Conveniently, all logical states fit entirely inside on-chip memories such as block RAM (BRAM). For applications such as SWAG, the concept behind its advantageous system behaviour is the replacement of the need for large indexed or hashed structures by offloading the design complexity to sorting, as it eliminates fragmentation originally caused by maintaining per-group information. FPGA-based high-throughput sorters are thoroughly studied in literature [saitoh2018high, zuluaga2016streaming, qiao2022topsort, song2016parallel], but in order to seamlessly support the median operator in our SWAG engine, innovative lightweight algorithmic modifications are introduced.
The contributions of Enthuse are as follows:
@@ II Enthuse – aggregation framework
The main idea behind Enthuse is that the pre-existing prefix-scan topology of a sparsity-removal circuit can be exploited and reused to also calculate and append aggregation information to the data, before filtering redundant tuples. The focus of this pipeline is high-throughput, as it is always able to process up to $P$ elements per cycle. It is inspired by the PRRA architecture [prra], where a rolling prefix scan is used to calculate the permutation indices for the permutation network (reverse butterfly) to achieve stream compaction (see figure 16, left).
@@ II-A Supported functionality
This design is a non-blocking building block working as a pipeline, i.e. it never asserts backpressure on the stream, as with the PRRA architecture. On the contrary, aggregation implies summarisation, hence the need to remove a newly generated intermediate sparsity with the compaction functionality of the permutation network [prra]. For most aggregation tasks, this will result in only the last tuple per group being considered useful, since it is repurposed to carry all the relevant accumulated information, while the remaining aforementioned “redundant” tuples may carry partial results.
@@ II-B Pipeline steps
Figure 1 introduces the architecture of Enthuse, a high-throughput pipeline for calculating a wide range of aggregation tasks including group-by. As an adaptable framework, it consists of a series of generalisable steps:
@@
the entities named $n$ are distributed inside the adapted prefix scan to calculate the permutation index of the tuples to achieve the round-robin effect by the reverse butterfly network, as with the PRRA [prra]. At the same time they also operate on the keykey fields, which later becomes the resultresult field containing the aggregation summary per group. For example, if the query function is to find the sum of all the keys for each of the group IDs, the entities $n$ not only perform a cumulative sum for the reverse butterfly indexindex, but also for the key sum per group, until the last entry of the same group is processed to carry the full sum of the group. An equality check on the group field with the previous entry ($group_{i}\stackrel{{\scriptstyle?}}{{=}}group_{i-1}$) is enough for knowing where not to propagate the previous result. Using sum as an example, the start of a new groupgroup would leave the current keykey intact instead of adding the result of the last tuple.
@@
finally, the reverse butterfly network permutes the batch of results to achieve the round-robin effect in parallel (see figure 16, left), and is inline with [prra]. The reverse butterfly network is used as an optimal-complexity switch, though other permutation networks can also be used. The lastlast flag is now renamed to a validvalid bit, as it now indicates that a port has a valid result.
@@ II-C Numerical example
Figure 3: High-throughput adaptable sliding window aggregation using the six-step pipeline of EnthuseSWAG.
@@
III EnthuseSWAG – Sliding window aggregation
The design of Enthuse from section II-B is adapted to perform sliding-window aggregation queries. In terms of functionality, the main difference from Enthuse is that the input is processed in a windowed context.
@@ III EnthuseSWAG – Sliding window aggregation
SWAG queries are usually of the form “calculate the median of the last WS elements every WA tuples” [geethakumari2023stream], where median can be replaced by any supported aggregate operator, WS is the window size, and WA is the distance between the sampled windows. Figure 4 shows an instance of the average operator, where on every 2 tuples (WA) we find the average key within the last 4 tuples (WS). In the literature, WS is also referred to as the window range, and WA is also referred to as the bulk size, slide or stride. In some contexts, there is a group ID associated with each key, thus, the engine can also provide per group information, as with Enthuse.
@@ III-A Supported functionality
The proposed SWAG engine supports the operators of the base pipeline (see section II-A), plus selection-based operators including the median. Based on the query classification from [Tangwongsan2018], the supported types are Sum-like, Max-like and Sketch-like.
In contrast to software algorithms that calculate SWAG incrementally [Tangwongsan2018], which implement the window eviction post calculation [tangwongsan2017low], here, there is no requirement for an inverse function. This is because each window is processed in its entirety, which contributes to the adaptability of the EnthuseSWAG for operators that cannot be calculated incrementally. Section VI elaborates on the applicability of Enthuse and EnthuseSWAG to other prominent aggregation workloads.
@@ III-B Design
The input data are seamlessly sorted on-the-fly using a small-scale sorter (see section III-D). As opposed to the pipeline of Enthuse, the input is not expected to be sorted, since the latency of doing so is shown to be minimal for the window sizes of interest. Figure 3 introduces the pipeline of EnthuseSWAG in steps, for data arriving in batches of $P$ tuples.
@@
a sorter is needed right before the aggregation engine to sort the window based on the groupgroup IDs, as well as the keys. Note that sorting by the keys as well is not always required, such as when the operator is commutative as with summation. Since 4K elements, for example, are considered moderately large as SWAG window sizes [geethakumari2023stream], the sorting can entirely be done on-the-fly instead of indexing larger memories. FPGA-based high-throughput sorters are thoroughly studied in literature [saitoh2018high, zuluaga2016streaming, qiao2022topsort, song2016parallel].
@@
the last two steps reduce back to a PRRA architecture [prra], as the remaining functionality only relates to the round-robin compaction. The final index calculation is done by the last prefix scan in the pipeline, while the first one is now responsible for all computation-related functionality including the other two cumulative sums.
@@ III-D Sorter with group-cardinality calculation
The ability to involve median calculation within a sorter depends on the sorter architecture [guo2001integrated]. Our proposed solution is based on a sorter that uses the linear sorter as a building block [fsorter]. The linear sorter is the hardware equivalent of insertion sort, and is able to insert a new value into an already sorted sublist [fpgasort]. Each sorting cell, as depicted in figure 6 (center), has a comparator that compares its tuple with the incoming one. According to the result of the comparison and the one of its neighbour (“left smaller”), it either adopts the new value, propagates the left value, or remains as is.
@@
The baseline sorter of our design is able to provide high-throughput (multiple tuples per cycle) while supporting moderately large window sizes. The open-source sorter of [fsorter] is specifically selected because it combines linear sorters with a merge tree (PMT [song2016parallel]) for high-throughput. A special feature of this combination is that it is able to work efficiently in many-leaf (input list) merge mode using the same hardware, and this also applies to our new logic for resource efficiency.
@@
An additional modification is needed to propagate the correct group cardinalities, as summarised in figure 8. The parallel merge tree (PMT) needs to sum the cardinalities calculated by the linear sorters. Each FLiMS merger [flimsj] of the PMT uses a network of compare-and-swap (CAS) units, which are sorters of 2-elements. The modification here is that each CAS unit uses the cond1 and cond2 conditions to determine how to update two incoming cardinalities every time they share the same group ID. The tuples carry two extra bits inside the mergers, one for denoting if the updating of their cardinalities has already happened, and one for their original source (lists A or B according to figure 8). Cond1 (to sum both cardinalities) is true whenever the cardinality update has not already happened to any of the 2 inputs, and the source is different. Cond2 (to adopt the other cardinality) is true whenever the cardinality update has happened to the other tuple, and the source is either different or the same. A similar arrangement is done for the selector stage units, as they are based on the same comparison, but they also remember any merged cardinalities from the elements of the last comparison. The latter is needed to support the case when a set of same group IDs is only coming from one of the lists A or B.
@@ IV Evaluation
This section evaluates the implementations of the group-by aggregate and SWAG engines as accelerators. All implementations are accompanied by a sorter/ cardinality counter of capacity $k$ of 128 elements, supporting arbitrary input sizes for the group-by engine by being applied recursively, and window sizes of up to $\textit{WS}=16K$ for EnthuseSWAG. The operator type is configured at run-time. The implemented engines support multiple operators also configurable at run-time, as summarised in table I.
@@ IV-A Experimental methodology
The engines are evaluated as co-processors in a Linux-running environment. They are implemented as AXI peripherals, and the codebase is mainly written in Verilog [^1] Source code: available soon at https://philippos.info/nca/software . The entities such as the compare-and-swap units for the sorter’s merger cells or the entities of the prefix scans are parameterisable, such as with custom data widths and functionality, and are written by hand. The structures that use those entities including the reverse butterfly network and the parallel merge tree in the sorter are generated by scripts. These are written in Python to easily infer such complex or repetitive entity topologies for arbitrary $P$ and $k$ (i.e. sorter size) values.
@@
The 2GB memory of the board is divided in two halves by the kernel; the kernel-mapped part, which is optimised for cached access to the CPU, and the non-kernel part, which is left for the FPGA. The target operating frequency of the AXI-related IPs is 250 MHz, with an internal frequency of 125 MHz for the adapted sorter and engines, complying with the sorter’s original design. This results in a 256-bit/cycle datapath for the engine from an 128-bit-wide HP port. This is relatively close to saturating the memory bandwidth of this platform [manev2019unexpected]. Figure 9 visualises this experimental setup.
@@
The data first pass through the high-throughput sorter, and its output is directed towards the rest of the aggregation engine. The engine is either Enthuse for group-by-aggregation or the corresponding part of EnthuseSWAG for sliding-windows. For the first, the sorter is the unmodified open-source IP from [fsorter], whereas the latter combines the group cardinality functionality from section III-D.
@@
Their performance is compared against highly-optimised software, including reference implementations from IBM for SWAG [ibm]. The custom operators remain relatively simple, as they rely heavily on well-regarded functions and structures from C++’s standard template library (STL). The baseline software uses std::set and std::map to hold and update per-group information (notably similar performance to their hash-based unordered counterparts, not shown), and std::nth_element() to calculate the medians. The GCC compilation flags are -O3 -march=native.
IV-B Experimental results – Enthuse, group-by aggregation